OpenStudy (anonymous):

Simplify this radical √-4 Does it matter it's a negative? Will it eventually become positive?

jimthompson5910 (jim_thompson5910):

\[\large \sqrt{-4} = \sqrt{-1*4}\] \[\large \sqrt{-4} = \sqrt{-1}*\sqrt{4}\] \[\large \sqrt{-4} = i*2\] \[\large \sqrt{-4} = 2i\]

jimthompson5910 (jim_thompson5910):

by definition, \[\large i = \sqrt{-1}\]

OpenStudy (anonymous):

wow that was super quick. :)

jimthompson5910 (jim_thompson5910):

glad to be of help

OpenStudy (anonymous):

what is i imaginary number

jimthompson5910 (jim_thompson5910):

yes, defined to be the square root of -1

jimthompson5910 (jim_thompson5910):

it's not a fair name since all numbers are imaginary really

OpenStudy (unklerhaukus):

are you confusing imaginary and complex @jim_thompson5910 ?

jimthompson5910 (jim_thompson5910):

no, i is an imaginary number

jimthompson5910 (jim_thompson5910):

something like 3+i is a complex number i is also complex as well

jimthompson5910 (jim_thompson5910):

since i = 0+1i

OpenStudy (anonymous):

True, so having a negative number does affect the outcome. I read it didn't because both negatives will become positive

OpenStudy (unklerhaukus):

\(i\) in not a complex number, \(i\) is a purely imaginary number

jimthompson5910 (jim_thompson5910):

not sure what you mean exactly newatthis

jimthompson5910 (jim_thompson5910):

i is in the complex number set

jimthompson5910 (jim_thompson5910):

i = 0+1i it's of the form a+bi where a = 0, b = 1

OpenStudy (anonymous):

I'm just a bit confused with negative √ simplified radical

OpenStudy (unklerhaukus):

i thought to be complex a and b were necessarily non-zero

jimthompson5910 (jim_thompson5910):

just keep in mind that x^2 is always positive or 0 so if x^2 = -4 for instance, then there are no real solutions but if you make the definition that i = sqrt(-1), then you can come up with 2 complex solutions

OpenStudy (anonymous):

Nope, as a matter of fact 2 is a complex number also. Complex space enclose both real and imaginary

jimthompson5910 (jim_thompson5910):

well yeah if you had the number 3, it's a real number, but it's also part of the complex number system so 3 is technically complex although we'd just say it's real

jimthompson5910 (jim_thompson5910):

yeah akorn103 has the right idea

OpenStudy (anonymous):

I think he just misunderstood you when you said that all numbers were imaginary and thought you were making a mathematical and not "philosophical" statement hah

OpenStudy (anonymous):

I'm lost but I appreciate the help guys.

jimthompson5910 (jim_thompson5910):

just follow the steps I outlined (write them in your notes if that helps) and keep in mind that \[\large i = \sqrt{-1}\]

OpenStudy (anonymous):

i is negative -1 so -1 is an imaginary number correct?

jimthompson5910 (jim_thompson5910):

i is the square root of -1

OpenStudy (anonymous):

Thank you guys your so sweet I appreciate this :)

OpenStudy (unklerhaukus):

ok i concede , complex numbers are a superset of real and imaginary numbers but its like calling a square a rectangle (its true, but misleading )

OpenStudy (anonymous):

ehh, it gets more important in linear algebra and beyond